OpenStudy (anonymous):

if the hypotenuse is 41 inches and one of the legs is 40 inches what is the length of the other leg in inches?

OpenStudy (anonymous):

Hey I need go soon

OpenStudy (anonymous):

But i will help you as much as possible

OpenStudy (anonymous):

A^2 + B^2 = C^2 you know this rule right??

OpenStudy (anonymous):

So if you just do.. 40^2 + b^2 = 41^2

OpenStudy (anonymous):

you end up with 1600 + b^2 = 1681

OpenStudy (anonymous):

Subtract 1600 on both side and end up with B^2 = 81

OpenStudy (anonymous):

which does make sense in this case because triangle inequality rule says " for any triangle, two sides added must be larger than the hypotenuse"

OpenStudy (anonymous):

40 + 9 = 49 and 49 is greater than 41. so it makes it a triangle. It is just a nice way to check if your triangle makes sense :). Leave me a pm if you have a question. I come back in 4 hours.